Each of 100 firms in a perfectly competitive industry has an AVC function AVC= 2Q

where Q represents the output produced. If the price of the product is Rs.100 find the total

quantity supplied in the market when a) total fixed cost is 10,000 and b) total fixed cost is

Rs.20,000.


1
Expert's answer
2021-02-22T15:12:05-0500

The average variable cost is AVC=2q.

The total variable cost will be equal to TVC=2q^2.

The marginal cost is equal to:

"MC=\dfrac{d TVC}{d q}\\[0.3cm]
MC=4q"

For a perfectly competitive firm, P=MC. Therefore, the supply curve for an individual firm is:

"P=4q\\[0.3cm]
q=0.25P"

There are 100 firms in the industry. Thus, the market supply curve is:

"Q=100(0.25P)\\[0.3cm]
Q=25P"

At a price of P=100, the total industry supply is:

"Q=25(100)\\[0.3cm]
Q=2500"

A change in the fixed cost will not affect the market supply.